What is the mass of
        (b) 0.5 mole of water molecules?

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Mole of water molecule = 0.5 mole
Molar mass of water molecules = 2 × 1 + 16 = 18 g
Mass of H2O = 18 × 0.5 = 9 g



Thanks